# Understanding the Application of Utility Theory in Robotics and Artificial Intelligence: A Survey

### v1

@@
Abstract — As a unifying concept in economics, game theory, and operations research, even in the Robotics and AI field, the utility is used to evaluate the level of individual needs, preferences, and interests. Especially for decision-making and learning in multi-agent/robot systems (MAS/MRS), a suitable utility model can guide agents in choosing reasonable strategies to achieve their current needs and learning to cooperate and organize their behaviors, optimizing the system’s utility, building stable and reliable relationships, and guaranteeing each group member’s sustainable development, similar to the human society. Although these systems’ complex, large-scale, and long-term behaviors are strongly determined by the fundamental characteristics of the underlying relationships, there has been less discussion on the theoretical aspects of mechanisms and the fields of applications in Robotics and AI. This paper introduces a utility-orient needs paradigm to describe and evaluate inter and outer relationships among agents’ interactions. Then, we survey existing literature in relevant fields to support it and propose several promising research directions along with some open problems deemed necessary for further investigations.
I INTRODUCTION
When people study, analyze, and design robotic or artificial intelligence (AI) systems, they would like to compare them with natural systems working. For example, many natural systems (e.g., brains, immune system, ecology, societies) are characterized by apparently complex behaviors that emerge as a result of often nonlinear spatiotemporal interactions among a large number of component systems at different levels of the organization [1]. From the single-robot perspective, cognitive developmental robotics (CDR) aims to provide a new understanding of how human higher cognitive functions develop by using a synthetic approach that developmentally constructs cognitive functions [2]. For multi-agent systems (MAS) [^1] In this paper, we use the term agent to mean autonomous intelligent entities, such as robot., they potentially share the properties of swarm intelligence [3] in practical applications (like search, rescue, mining, map construction, exploration, etc.), representing the collective behavior of distributed and self-organized systems. There are a lot of research fields related to it, especially building so-called artificial social systems, such as drones and self-driving cars. Specifically, in multi-robot systems (MRS), each robot can internally estimate the utility of executing an action. And robots’ utility estimates will be inexact for several reasons, including mission cost, strategy rewards, sensor noise, general uncertainty, and environmental change [4, 5].
Figure 1: Rewards (Cumulative Utilities) Driven Behaviors
Furthermore, in MAS cooperation, trust is an essential component for describing the agent’s purposes. In other words, there are complex relationships between utility, transparency, and trust, which depend on individual purposes [6]. Especially considering humans and robots collaborated as a team, recognizing the utility of robots will increase interest in group understanding and improve Human-Robot Interaction (HRI) performance in tasks [7]. In this paper, we introduce a utility-orient needs paradigm and survey existing work on robotics and AI through the lens of utility. We organize this review by the relationships between single-agent, multi-agent, trust among agents, and HRI to underscore the effects of utility on interaction outcomes.
I-A Scope and Contributions
@@
One of the significant challenges in designing robotic and AI systems is how to appropriately describe the relationships of their interconnected elements and natures. It is natural to introduce a unifying concept expressing various components and evaluating diverse factors on the same ground. As a results, Utility Theory is a wide cross-area of research that lies at the intersection of economics, operations research, robotics, AI, cognitive science, and psychology. To the best of our knowledge, this survey is the first to formalize a comprehensive model of the utility paradigm in agents’ interaction from agents’ needs perspective and situate prior and ongoing research regarding how the utility can influence inter and outer relationships among agents and affect systems’ performance. The main contributions of this paper are listed as follow:
  • We introduce a utility paradigm based on agents’ needs to describe and evaluate inter and outer relationships among agents’ interaction, integrating and summarizing insights from robotics and AI studies.
  • We review existing literature in relevant fields that support the proposed utility-orient needs paradigm of agents’ interaction.
  • We propose several open problems for future research based on utility theory, which will support researchers in building robust, reliable, and efficient robotic and AI systems, especially in the MAS and HRI domain.
@@
Fig. 2 outlines the vital relationships of the utility-orient needs paradigm based on agents’ motivations among single-agent, MAS, and HRI, which affect systems’ abilities to meet the requirements of specific tasks and adapt to human needs and uncertain environments. We survey how the utility-orient relationships affect the performance of different systems, especially evaluating the trust level among agents through various types of utilities and integrating human needs and agents’ utilities to build a harmonious team.
II Synopsis of Affordance Formalisms
@@
This section summarises the evolution of formalisms that use the concept of utilities to improve agents’ performance. This evolution can be classified into two main stages. Mathematical conceptualizations characterize the first stage as extensions of theories from various fields. The second stage corresponds to formalisms that focus on optimizing systems’ capabilities by unifying different concepts based on Utility Theory.
II-A Utility Theory

### v3

@@
Abstract
Cognitive modeling, which explores the essence of cognition, including motivation, emotion, and perception, has been widely applied in the artificial intelligence (AI) agent domains, such as robotics. From the computational perspective, various cognitive functionalities have been developed through utility theory to provide a detailed and process-based understanding for specifying corresponding computational models of representations, mechanisms, and processes. Especially for decision-making and learning in multi-agent/robot systems (MAS/MRS), a suitable cognitive model can guide agents in choosing reasonable strategies to achieve their current needs and learning to cooperate and organize their behaviors, optimizing the system’s utility, building stable and reliable relationships, and guaranteeing each group member’s sustainable development, similar to the human society. This survey examines existing robotic systems for developmental cognitive models in the context of utility theory. We discuss the evolution of cognitive modeling in robotics from behavior-based robotics (BBR) and cognitive architectures to the properties of value systems in robots, such as the studies on motivations as artificial value systems, and the utility theory based cognitive modeling for generating and updating strategies in robotic interactions. Then, we examine the extent to which existing value systems support the application of robotics from an AI agent cognitive modeling perspective, including single-agent and multi-agent systems, trust among agents, and human-robot interaction. Finally, we survey the existing literature of current value systems in relevant fields and propose several promising research directions, along with some open problems that we deem necessary for further investigation.
keywords:
[…]
{graphicalabstract}
{highlights}
Robots’ Motivation Mechanism Definition and Modeling
Robots’ Value System Cognitive Modeling through Utility Theory
Artifical Social System integrating with Human Society
1 Introduction
When people study, analyze, and design artificial intelligence (AI) agent or robotic systems, they would like to compare them with natural systems working. For a living organism, cognition refers to its capacity to process perceptual information and thereby manipulate its behaviors, which reflects the level of complexity and intelligence of the agent [173]. For example, many natural systems (e.g., brains, immune system, ecology, societies) are characterized by apparently complex behaviors that emerge as a result of often nonlinear spatiotemporal interactions among a large number of component systems at different levels of the organization [142]. Moreover, the cognitive capabilities of humans are generally regarded as the manifestation of self-awareness, perception, learning, knowledge, reasoning, planning, and decision-making [26]. The related research primarily focused on enhancing the capabilities of artificial systems and deepening our understanding of the fundamental roots of intelligence in natural systems [146, 269], such as artificial development [225], epigenetic robotics [182], development robotics [53, 151], and cognitive developmental robotics (CDR) [14]. Specifically, artificial or mental development refers to a brain-like natural or artificial system that improves its capabilities under the control of its species-specific developmental program [269]. Developmental and epigenetic robotics studies the mechanisms, architectures, and constraints that allow lifelong and open-ended learning of new skills and knowledge in embodied machines [151, 182]. Furthermore, cognitive developmental robotics aims to provide a new understanding of how human higher cognitive functions develop by using a synthetic approach that developmentally constructs cognitive functions [15].
@@ 1 Introduction
The history of cognitive modeling in robotics can be traced back to behavior-based robotics (BBR) [13], which is based on a biologically inspired philosophy supporting parallel, decentralized architectures and allowing some freedom of interpretation. BBR emphasizes the “brains" of the robots rather than their “bodies", which moves the details of robot sensors and actuators firmly into the background. It tries to understand what behaviors we should expect robots to exhibit and which computational mechanisms can achieve them. In behavior-based systems, the robot controller consists of a collection of behaviors, each of which achieves and/or maintains a specific goal, such as avoiding collision, finding targets, and wandering [167]. In the multi-robot systems (MRS) and swarm robotics domain, behavior-based modeling focuses on developing generic mechanisms that result in coordinated group behavior, either implicitly or explicitly [70], such as formation control [19] and cooperative architecture design [203]. However, BBR mainly focuses on simple behaviors and rule-based design, and it is hard to achieve complex behaviors and build sophisticated relationships among agents (like trust, respect, and ethic), especially in human-robot interaction (HRI).
With the development of AI and cognitive science, cognitive architectures are increasingly incorporated into autonomous robot design, giving rise to new disciplines, such as cognitive robotics [52], and promoting the application of cognitive modeling in robotics. Specifically, cognitive modeling deals with implanting human-like cognition, albeit reduced complexity, in autonomous robots with the hope that the cognitive capacities, as they work in humans and other biological creatures, will also assist the new generation of robots to be coherent, self-motivated, social, persistent in behavior, and aware of themselves and their environments [26]. It combines the neuroanatomy, cognitive neuroscience, neuroeconomics, cognitive psychology, developmental studies, linguistics, and AI with robotics. The majority of research on cognitive modeling in robots focuses on developing models of discrete cognitive abilities, such as attention, learning, value systems, reasoning, and decision-making. The majority of research on cognitive modeling in robots focuses on developing models of discrete cognitive abilities, such as attention, learning, value systems, reasoning, and decision-making. In particular, studies on developing clear metrics and benchmarks for the different aspects of HRI aim to endow robots with advanced cognitive and communicative abilities [10].
From multi-agent systems (MAS) perspective, cognitive models can provide a substantial way to better probe multi-agent issues by taking into account the essential characteristics of cognitive agents and their various capacities [244]. Furthermore, cognitive modeling presents a more realistic basis for understanding multi-agent interaction by embodying realistic constraints (like agents’ innate-values and needs), capabilities, and tendencies of individual agents in interacting with their environments, including physical and social environments [242]. Agents potentially share the properties of swarm intelligence [27] in practical applications (like search, rescue, mining, map construction, exploration, etc.), representing the collective behavior of distributed and self-organized systems. There are a lot of research fields related to it, especially building so-called artificial social systems, such as drones and self-driving cars. Specifically, in multi-robot systems (MRS), each robot can internally estimate the utility of executing an action. And robots’ utility estimates will be inexact for several reasons, including mission cost, strategy rewards, sensor noise, general uncertainty, and environmental change [95, 281]. Furthermore, in MAS cooperation, trust is an essential component for describing the agent’s purposes. In other words, there are complex relationships between utility, transparency, and trust, which depend on individual purposes [275]. Especially considering humans and robots collaborated as a team, recognizing the utility of robots will increase interest in group understanding and improve Human-Robot Interaction (HRI) performance in tasks [210].
However, bringing all of these capabilities, such as self-awareness, perception, knowledge, reasoning, planning, decision-making, and learning, together in a single artificial agent (a robot) remains one of the grand challenges, let alone multi-agent systems and human-robot interaction. One of the significant challenges in designing AI and robotics systems is how to formalize and quantify motivations [268, 267, 14, 252]. The problem of producing structures that address the compositional construction of useful knowledge representations, relating the innate drives and needs of an intelligent system to the characteristics of the environment in which it finds itself, is of paramount importance [220]. Another challenges is how to appropriately describe the relationships of their interconnected elements and natures. Related issues include developing coordination strategies (that enable groups of agents to solve problems together effectively), negotiation mechanisms (that facilitate the interaction of a set of agents), conflict detection and resolution strategies, and other mechanisms whereby agents can contribute to overall system effectiveness while maintaining autonomy.
@@
Generally speaking, establishing goal/sub-goal hierarchies in agents’ motivational systems for open-ended learning situations and modeling their value systems based on utility theory is an open research problem, especially when their state spaces are continuous and ambiguous [219, 23]. Furthermore, in social scenarios, the interaction dynamics are typically unknown and unpredictable, so robots working in these environments must have appropriate decision-making capabilities to select their actions and successfully fulfill the intended tasks autonomously [160]. Fig. 1 outlines the vital relationships of the utility-orient needs paradigm based on agents’ motivations among single-agent, MAS, and HRI, which affect systems’ abilities to meet the requirements of specific tasks and adapt to human needs and uncertain environments. It presents how the utility-orient relationships affect the performance of different systems, especially evaluating the trust level among agents through various types of utilities and integrating human needs and agents’ utilities to build a harmonious team [277].
Figure 2: The illustration of relationships between the cognition evolution and robot development [120].
Following the classical expectancy-value theory [272], the utility theory is regarded as a paradigm for describing the goals and motivations of an AI agent. A utility model serves as a function that maps the utility to a goal for each point in the state space, allowing goals to be reached by following paths of increasing utility within the state space. Previously, K. Merrick [173] discussed the value systems for developmental cognitive robotics. More recently, M. Maroto-G´omez [160] provided a comprehensive overview of control architectures for autonomous and social robots. However, a deeper analysis of the current state of utility-based cognitive modeling in robotics, including assessing its evolution over the years and framing its challenges and future goals, is necessary. Therefore, we propose this contribution, which fills a gap in the literature by providing a survey of utility theory based cognitive modeling in robotics.
The remainder of this paper is organized as follows: Section 2 presents the foundations of cognitive modeling in robotics, including BBR, cognitive architectures, and utility-based models in cognition. Sections 3 and 4 review the utility theory based cognitive modeling of single-agent and multi-agent systems in the theoretical foundation, fundamental AI architectures, and autonomous and social robots by area of application. Section 5 analyzes the categories of trust in robotics based on cognitive modeling and surveys the impressive body of work from recent years. Section 6 examines how utility-based value systems integrate into human-robot interaction and have been evaluated by the research community. Section 7 proposes several open questions for future research in related fields. We conclude in Section 8 with a discussion of our survey and provide our own experience as researchers.
2 Foundations of Cognitive Modeling in Robotics
@@
Cognitive modeling in robots involves multiply disciplines, such as robotics, AI, computer science, cognitive science, neuroscience, neuroeconomics [96], biology, control theory, decision making theory [83, 180], and psychology. This section presents the methodology related to the evolution of cognitive modeling from traditional behavior-based modeling and cognitive architectures to utilizing the concept of utility to improve cognitive robots’ performance. This evolution can be classified into two main stages. Behavior-based robotics characterize the first stage as extensions of theories from various fields. The second stage corresponds to cognitive robotics that focus on optimizing systems’ capabilities by unifying different concepts based on utility theory. Fig. 2 illustrates relationships between cognition evolution and robot development.
2.1 Behavior-based Robotics
Figure 3: The illustration of the information flow in classical AI (left panel) and in BBR (right panel).
2.1.1 Features of Behavior-based Robots
To achieve human-level intelligence, researchers have been working for about a century. Many efforts in robotics research have been inspired by simple biological organisms, with the aim of understanding and implementing basic survival-related behaviors in robots before progressing to more advanced behaviors involving, for example, high-level reasoning [261]. In the classical AI framework, the robot perceives the environment through sensors and builds a world model to predict the effects of various actions within it, enabling reasoning before executing behaviors in the real world. However, the procedure is very different from the distributed computation of the biological brains, and it is not the function of most biological organisms. Furthermore, Rodney Brooks argued that building world models and reasoning using explicit symbolic representational knowledge at best was an impediment to timely robotic response and at worst actually led robotics researchers in the wrong direction [45].
Distinct from traditional AI, behavior-based robotics (BRR) provides a paradigm for exhibiting complex behaviors with little internal state, modeling its immediate environment for adaptation, and gradually correcting robots’ actions via sensory-motor links [13]. In BBR, intelligent behavior can be defined as the ability to survive and to strive to reach other goals in an unstructured environment [^1] An unstructured environment is one that changes rapidly and unexpectedly, making it impossible to rely on predefined maps, which holds true for most real-world environments.. Fig. 3 illustrates the information flow of classical AI and the intelligent behaviors built into BRR in a bottom-up fashion, starting with simple behaviors executed simultaneously in a given robotic brain and providing suggestions for which actions the robot should take.
BBR is a branch of robotics that bridges AI, engineering, and cognitive science, providing approaches to embedded system control across research and practical applications. It aims to develop methods for controlling artificial systems, ranging from physical robots to simulated ones and other autonomous software agents, and to use robotics to model and understand biological systems more fully, typically, animals ranging from insects to humans [166]. Behavior-based systems employ a collection of concurrently executing behaviors, processes connecting between sensors and effectors. An important property of behavior-based systems is its ability to contain state, construct, and utilize distributed representations simultaneously [189].
2.1.2 System Organization
Figure 4: The illustration of the behavioral module in BRR.
In behavior-based systems, the robot controller consists of a set of behaviors that achieve and/or maintain specific goals. Each behavior is regarded as a processing element that serves as a control law and can be implemented in software or hardware. Moreover, it can receive information from the robot’s sensors, such as cameras and ultrasound, infrared, or tactile sensors, and/or from other behaviors, and send results to the robot’s effectors, like wheels, grippers, arms, or speech, and/or to other behaviors in the system [166]. From a control theory perspective, a structured network of interacting behaviors governs a behavior-based robot.
Different behavior-based robots have some common basic features, such as survival and safety. As a top priority, behavior-based robots need to provide the most basic behaviors, such as obstacle avoidance and battery charging, to ensure normal function in an unstructured environment. Then, by generating several behaviors in operation simultaneously, the final action taken by the robot corresponded either to an optimal choice among the suggestions from the various behaviors or to an average action formed from the suggested actions. To achieve autonomous robot interaction without direct human supervision, BRR introduces the critical concept of situatedness: Behavior-based robots do not build complex, abstract world models but adapt to real-world situations. Their behavior is reactive, with direct coupling between sensors and actuators. Moreover, their internal states usually reflect motivations stored in short-term memory, which generate current goals or long-term achievement.
Figure 5: The illustration of the simple three-layered robot architecture [45].
Generally speaking, a set of basic behaviors is regarded as a brain organ that organizes the actions of the behavior-based robot, termed the behavioral repertoire. The process involves two steps: defining individual behaviors and selecting actions. Especially in any robot intended for complex applications, the behavioral selection system for the complex decision-making process should balance various factors and is as important as the individual behaviors themselves [189].
The key challenge for behavior-based systems is coordinating multiple behaviors. To simplify the systems, most implementations use a built-in or fixed-priority ordering of behaviors. Considering computation efficiency and ease of analysis, common approaches select a behavior by computing a function of the behavior’s activation levels, such as voting or activation spreading [154, 204]. Behavior-based systems are related to the works of reactive robotics, particularly the subsumption architecture [43]. Through embedding the robot’s controller with a collection of preprogrammed parallel condition–action rules or reflexes in a minimal internal state, it can achieve rapid real-time responses [4, 47].
Specifically, task-achieving behaviors in the subsumption architecture are represented as separate layers. Individual layers work on individual goals concurrently and asynchronously. At the lowest level, each behavior is represented by a behavioral module. Where each behavioral module performs an action and is responsible for its own perception of the world, and there is no global memory, bus, or clock. (Fig. 4) that encapsulates a specific behavioral transformation function. Stimulus or response signals can be suppressed or inhibited by other active behaviors. A reset input is also used to restore the behavior to its initial state. A priority hierarchy fixes the topology of behavioral layers. The lower levels in the architecture are unaware of the higher levels, which provides the basis for incremental design. Higher-level competencies are added on top of an already functioning control system without modifying the lower levels. Fig. 5 shows a simple mobile robot with three behavioral layers.
To summarize, the key aspects for the design of subsumption-style robots in BRR are situatedness and embodiment. Situatedness refers to the robot’s ability to sense its current surroundings and avoid using abstract representations, while embodiment insists that robots be physical creatures and thus experience the world directly rather than through simulation [45, 13]. To some extent, BBR promotes the integration of robotics and cognitive science, and is a traditional embedded AI system [290] considered an example of weak artificial intelligence.
Table 1: Competency Areas and System Types of Typical Cognitive Architectures
Cognitive Architectures Perception Attention Motivation Reasoning Learning Actuation Interaction Creativity Type*
Soar: [134] ATC-R: [11] CLARION: [245] NARS: [265] LIDA: [77] SiMA: [228] Sigma: [211] CogPrime: [98] ARCADIA: [42] STAR: [256] CELTS: [78] PRODIGY: [75] MusiCog: [168] ATLANTIS: [94] CoSy: [199] 3T: [81] Leabra: [194] SASE: [270] $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ - $\bigstar$ $\bigstar$ $\bigstar$ - - $\bigstar$ - - - - - - - $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ - - $\bigstar$ $\bigstar$ $\bigstar$ - $\bigstar$ - $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ - $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ - $\bigstar$ - $\bigstar$ - $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ - $\bigstar$ $\bigstar$ $\bigstar$ $\bigstar$ - - $\bigstar$ $\bigstar$ - - $\bigstar$ $\bigstar$ - $\bigstar$ - - $\bigstar$ - - - - - - - - - $\bigstar$ - - - - - HA HA HA HA HA HA HA HA HA HA HA SA SA HA HA HA EA EA
    Note: Symbolic Architectures: SA; Emergent Architectures: EA; Hybrid Architectures: HA.
2.2 Cognitive Architectures
Instead of studying individual low-level or simple behaviors to model robots’ interactions in BBS, cognitive architectures stand for more general and abstract aspects, such as perception, attention, motivation, reasoning, learning, and creativity, to embody AI agents as creatures. They are part of Artificial General Intelligence (AGI) research. The goal of the research in cognitive architectures is to model the human mind, eventually enabling us to build human-level artificial intelligence [126]. Existing cognitive architectures have explored four perspectives of AI: mimicking human thinking styles, rational thinking, mimicking human behaviors, and rational actions [222], which provide evidence for particular mechanisms succeeding in generating intelligent behaviors. A cognitive architecture reflects three aspects of a cognitive agent [137], which are constant over time and across different application domains:
  • Having the short-term and long-term memories to store the agent’s beliefs, goals, experience, and knowledge;
  • The components and their organization represent larger-scale mental structures in memories;
  • The functional processes of corresponding structures involve actuators and learning mechanisms.
Specifically, a typical cognitive architecture usually contains several critical components, such as a perceptron for processing sensory inputs and translating them into meaningful information, an attention mechanism for selecting relevant stimuli, working memories for temporarily storing and manipulating information to perform tasks, long-term memories for storing and retrieving information through learning and experience, reasoning and decision making for processing information and making inferences to reach decisions, and action modules for executing motor actions that are necessary to achieve goals. Fig. 6 illustrates a cognitive architecture with the autonomy and learning capacity to achieve lifelong open-ended learning autonomy (LOLA) [^2] Lifelong open-ended learning autonomy emphasizes autonomous goal discovery, inherently tied to challenges like competence acquisition, curriculum learning, and adaptation in non-stationary environments [236, 217]..
Figure 6: The illustration of a cognitive architecture reaching the autonomy and learning capacity necessary to achieve LOLA [218].
2.2.1 Criteria of Cognitive Architectures
Due to the absence of a clear definition and a general theory of cognition, most cognitive architectures rely on particular premises and assumptions, making comparisons and evaluations of progress across architectures difficult [126]. Some researchers have argued that current models of cognitive architecture do not adequately reflect the complexity of human cognition, thereby limiting machine intelligence’s ability to replicate human decisions. Furthermore, they have not provided a theoretical framework for explaining how the mind’s cognitive processes interrelate. Although the current cognitive architecture models can explain how individual cognitive processes work, they don’t provide a comprehensive understanding of how these processes interact and shape human behavior.
To develop criteria for cognitive architecture, some researchers have done pioneering work. [186, 187] proposed the functional criteria, which include flexible behavior, real-time operation, rationality, a large knowledge base, learning, development, linguistic abilities, self-awareness, and brain realization. [243] extends the criteria to broader domains, including ecological, cognitive, bioevolutionary realism, adaptation, modularity, routineness, and synergistic interaction. Furthermore, some researchers proposed more practical solutions regarding it as a set of competencies and behaviors demonstrated by the system, such as perception, memory, attention, actuation, reasoning, learning, social interaction, planning, motivation, emotion, creativity, etc [2]. Table. 1 presents typical cognitive architectures and their competency areas.
2.2.2 Categories of Cognitive Architectures
Based on previous literature reviews [185, 126, 99], cognitive architectures can be regarded as control architectures and classified into three categories: Symbolic, Emergent, and Hybrid. Fig. 7 illustrates the classification of cognitive architectures and their corresponding control architectures.
Symbolic Architectures:
Symbolic architectures are closely related to cognitivism, which views the human mind as processing information through symbolic representations and logical rules. They utilize formal languages to represent information and rules guiding decision-making. Their hierarchical structure can set high-level goals to organize task planning and execution. Like the deliberative architectures, symbolic architectures also focus on information processing based on symbolic knowledge. For example, robots can use symbolic information and logical rules to achieve high-level abstraction and long-term planning, such as route planning in unstructured environments, natural language processing, logical inference, and real-time decision-making [135]. Moreover, the properties of symbolic architectures make them particularly suitable for formal verification and safety-aware planning, thereby contributing to safety, verification, and explainability in robotic systems [152, 33].
Emergent Architectures:
Unlike symbolic architectures based on symbolic representations or logical rules, emergent architectures build on complex behaviors and capabilities emerging from the interaction of multiple simple elements. On the other hand, reactive architectures are built on predefined sets of rules and behaviors specific to particular situations. Although they share features with reactive architectures in perception-based decision-making and adaptation, emergent architectures use neural networks and other machine learning techniques to generate behaviors, often learning and adapting based on experience and environmental feedback. Moreover, the enactive interpretation attribute causes them to adjust their behavior as they explore new situations or receive new data, which might sabotage the previously learned model. Recent studies unify views of perception, action, and learning based on the free-energy principle, serving as architectural principles for embodied agents [88, 207].
Hybrid Architectures:
As we discussed above, symbolic architectures are good at well-defined problems, such as planning and reasoning. Their knowledge base is based on human interpretation and needs to be generated from initial states, which makes them hard to adapt to dynamic environments. Although emergent architectures are highly adaptable and exhibit enactive interpretation, they might degrade model performance by using sub-symbolic knowledge and exploring new data. To tackle those issues, hybrid architectures attempt to combine two frameworks into a single structure. They usually have a symbolic planner in charge of long-term planning, and several emergent modules that provide actions with greater adaptability. For example, the neuro-symbolic integration incorporates symbolic reasoning and planners with neural perception and learning modules, involving translating neural outputs to symbols and tightly coupled systems to interleave learning and reasoning [33]. Moreover, some foundation models, such as Large Language Models (LLMs) and multimodal Vision-Language Models (VLLs), accelerate the integration of perception, reasoning, and low-level skills by serving as high-level planners, communicators, and knowledge sources, generating action sequences or symbolic plans [5, 110, 264, 179, 116].
Figure 7: Illustration of the classification of cognitive architectures and their corresponding control architectures.
2.2.3 Cognitive Robotics
Cognitive architectures play an essential role in the development of intelligent autonomous robots, including planning and control, decision-making, learning, and adaptation to dynamic environments, which lay the foundation for cognitive robotics. Specifically, developmental and cognitive robotics studies the mechanisms, architectures, and constraints that enable lifelong, open-ended learning of new skills and knowledge in embodied machines. They aim to establish cognitive functions through a synthetic approach to better understand higher cognitive behaviors in humans [14], employing knowledge representation, reasoning, and learning techniques to support autonomous robots in dynamic and unstructured environments [141]. In the early stage of cognitive robotics studies, various robots, such as mobile robots, robot arms, and humanoids, were used to investigate higher-order cognitive capabilities that mimic the functionalities of the human brain, including its internal structure, infrastructure, and social structures.
Social Cognition:
Social cognition refers to the ability to recognize and control the self in relation to others and to perceive and apply social signals in the interaction. It extends cognitive robotics into a more general and sophisticated domain that not only considers multi-robot exchange of information and knowledge and the acquisition of new skills from others, but also involves building reliable and sustainable relationships with humans in the long-term development. For example, research on socially interactive robots [84, 36] is interested in very high-level social interaction conducted by robotic agents, focusing on the body of research dedicated to designing task non-specific social cognition in robotic agents as part of fulfilling the goal of designing human-like cognition in robotic systems. Therefore, “sociable robot" research on social cognition development in autonomous systems primarily focuses on mechanisms that enable robots to interact with humans in two main categories, including Joint Attention which refers to the ability to intentionally attend to an object/region of mutual interest [41, 46, 37, 184, 109], and Social Imitation which includes sensori-motor skills for generating different types of social cues and task sequence learning during development [40, 130, 235, 113, 50, 51].
Research Directions:
The core challenge of cognitive robotics is the gap between the robot’s knowledge and its reasoning capabilities [254], which makes automatic acquisition and deployment more difficult. Especially, let robots share common-sense knowledge (CSK) with humans in everyday tasks, such as knowledge about human desires, physics, causality, emotion, properties, and relationships [105]. This knowledge relates to the concept of Intuitive Psychology [136], which holds that an agent understands other agents with a mental state similar to their own, and can express and interpret this to achieve mutual understanding of their intentions and goals. One possible solution is Physical Embodiment, which means that the agent’s physical body specifies the constraints on the interaction between the agent and its environment that generate the rich contents of its process or consequences [44, 16, 15, 257, 131, 14]. Inspired by the concept, researchers build artificial systems capable of acquiring motor and cognitive capabilities through interaction with the environment, in line with human development [14]. Another approach gradually implemented in cognitive robotics is predictive processing (PP), a top-down approach that aims to integrate perception, cognition, and action as a single inference process [61], grounded in the free-energy principle [49] and frameworks such as predictive coding, active inference, and perceptual inference.
Furthermore, the development of machine learning techniques has led to a range of learning paradigms, including imitation learning, transfer learning, representation learning, reinforcement learning, and curriculum learning [250]. It provides systematic approaches, ranging from mastering lower-level basic skills, such as trajectory planning and multi-DOF (Degrees of Freedom) control [288], to more sophisticated systems, including modules for visual attention, speech recognition, and the integration of visual and linguistic inputs for instructing robots to grasp everyday objects [241]. As various techniques, such as AI, quantum computing, wireless sensors, and cloud computing, have rapidly developed, trends in cognitive robotics are presented below [251]:
  • Synthesizing computational models for high-order cognitive skills through neuroscience and behavioural psychology in AI agents;
  • Utilizing motivations or goal-directed mechanisms to balance exploration and exploitation in task spaces;
  • Using predictive coding mechanisms to synthesise higher-order cognitive behaviours;
  • Social behaviours studies in robotic swarms;
  • Large implementation of networked and cloud robotics and cyber-physical systems in higher-order robotic research, like social robots.
2.3 Utility Theory
2.3.1 Basic Concepts